at what displacement potential ewnergy is one 4th  of its maximum value?
 

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
 
Since it isn’t mentioned where but the question is likely to be of SHM so,

Let the amplitude is A
Potential Energy will be maximum when displacement is maximum from the mean position i.e A.
Potential Energy when x=A, U=kA2/2

ATQ,
kA2/8 = kx2/2
=>x=A/2.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)